# Multi-agent Adaptive Mechanism Design

### v1

@@ 2. Problem Formulation
  • reports a fixed permutation of her observation, i.e. for each player $i$ there exists a bijection $\pi_{i}:\mathcal{Y}\!\to\!\mathcal{Y}$ such that $Z_{it}=\pi_{i}(X_{it})$ .
Proposition 2.1 is derived from Blackwell’s informativeness theorem (blackwell1953equivalent). Each round, the true label, observations, reports, and decisions form a Markov chain: $Y_{t}\to\boldsymbol{X}_{t}\to\boldsymbol{Z}_{t}\to A_{t}$. An intuition is that optimal decision-making requires maximal information from upstream. Due to the data processing inequality (cover1999elements), information never increases going downstream; therefore, the best approach is to preserve as much information as possible at each link. Truthful reporting preserves full information at the link $\boldsymbol{X}_{t}\to\boldsymbol{Z}_{t}$, which allows for optimal subsequent decisions. Any lies from players erode information. Lazy behavior also produces less information than observing. Aside from truthfulness, an alternative case that preserves full information is when players use a permutation reporting strategy. However, such a case is unrealistic in practical settings, as the principal would need to know each player’s permutation rule to reverse the encoding and uncover the true observation. Therefore, this proposition essentially shows that eliciting truthfulness is the practical way to achieve maximal performance.
Truthfulness is necessary not only for data quality but also cost-optimality. Our adaptive mechanism relies on the idea of gradually learning players’ skills across tasks to lower costs in later rounds. Truthful reporting provides faithful data that allows accurate estimates of skills.
@@ 3. Mechanism Design without Common Knowledge
In our work, a central relaxation of modeling is that we don’t assume prior distribution of labels $p_{Y}$ or players’ skills $p_{i}$ are known by players or the principal. The principal’s central focus is to maintain truthfulness with unknown or inaccurate estimation of that knowledge. The absence of the so-called common knowledge assumption is a major distinction of our model from the classical mechanism design literature. We would follow Wilson’s critique, which appeals for relaxing the assumption of that player behaviors are common knowledge. In this section, we focus on (distributionally) robust mechanisms, which aim to incentivize truthful behavior under knowledge ambiguity. The notion of (distributionally) robust mechanisms is not new, and has been studied in (bergemann2005robust; koccyiugit2020distributionally) in general settings. Nonetheless, the mechanism in our section is specifically tailored to the estimation process of unknown player information, with the purpose of being applied in each stage of adaptive mechanisms.
@@ 3.1. Optimal Single-round Mechanism Design
We begin with the analysis of optimal mechanism design with known $p_{Y}$ and $p_{i}$ within a single round. When there are no true labels available, we apply the principles of peer prediction, which is to use other players’ report to verify a focal player’s report. The delicacy lies in the careful design of the reward mechanism to ensure that truthfulness is a Nash Equilibrium.
We start with the two-player mechanism. With a focal player $i$ and a reference player $j$, the optimal 2-player mechanism design problem could be formulated as a linear programming problem. The objective is to minimize expected reward to players, and the constraints are the desired properties of the mechanism, including truthfulness. (we hide subscript $t$ for simplicity.)
@@
The expectation is taken under the joint probability law $p_{\boldsymbol{X}}$ induced by $p_{Y},p_{i},$ and $p_{j}$. The first constraint is the individual rationality property, meaning when the player obtains observations and then reports honestly, she would get a non-negative reward. The second one states the truthfulness property, where the player receives a non-positive reward when she lies. The final constraint implements the no-free-lunch property, meaning the player cannot get a positive utility when she is lazy. We introduce the individual rationality and the no-free-lunch constraint to prevent arbitrary decrease of the objective function by applying an affine transformation to the reward mechanisms. (For risk-neutral players, affine transformations on reward do not affect utility ordering and strategic behavior.)
@@
Suppose there are two types of images $\mathcal{Y}=\{\texttt{Cat},\texttt{Tiger}\}$, abbreviated with $C$ and $T$ respectively. We further assume that the prior distribution of the image types is balanced, i.e., $p_{Y}(C)=p_{Y}(T)=0.5$. For each image with an unknown true label $Y\in\mathcal{Y}$, the principal would like to let two players $1,2$ individually observe it, and truthfully report their observations $X_{1},X_{2}$ to label that image. Assume that both players are $90\%$ accurate: $p_{i}(C\mid C)=p_{i}(T\mid T)=0.9$, and $p_{i}(T\mid C)=p_{i}(C\mid T)=0.1$.
The principal designs the reward mechanism as follows: both players receive $1$ reward if their reports $Z_{1},Z_{2}$ agree with each other, and receive 1-1 otherwise, i.e. $R_{\{1,2\}}(Z_{1},Z_{2})=2\cdot\mathbf{1}_{[Z_{1}=Z_{2}]}-1$. and assume that observation incurs cost $c=0.1$. Now we assume that player $2$ observes and report honestly, and analyze the incentive of player $1$.
Suppose player $1$ observes Cat, Bayes’ formula gives that $\mathbb{P}(X_{2}=C\mid X_{1}=C)=0.82$ and $\mathbb{P}(X_{2}=C\mid X_{1}=C)=0.18$. On the other hand, if player $1$ does not pay the effort to toss the coin, then her Bayesian belief on player $2$’s observation (and report) is $P(X_{2}=C)=P(X_{2}=T)=0.5$. She can then work out expected reward under truthful, lying, and lazy strategies: truthful ($0.54$) ¿ lazy (0) ¿ lying (0.74-0.74). Hence truthful behavior is desired. In fact, this simple mechanism is a feasible solution to Eq.(1). Intuitively, after the focal player’s observation, her posterior probability on the other player’s observing the same label is higher than observing a different label. Therefore, it is preferable to report whatever you observe in the first place. Such a mechanism is called peer prediction (miller2005eliciting), originated the fact that rational players always tries to predict their peers’ observations before action.
Define the belief matrix $\mathbf{B}$ where $\mathbf{B}_{xx^{\prime}}=\mathbb{P}(X_{j}=x^{\prime}\mid X_{i}=x)$, and reward matrix $\mathbf{R}$ where $\mathbf{R}_{xx^{\prime}}=R_{i}(x,x^{\prime})$. We also let $\mathbf{d}$ a column vector of the prior distribution of $j$’s observation: $\mathbf{d}_{x}=\mathbb{P}(X_{j}=x)$. Then we can reformulate (1) into the following equivalent problem:
@@
Note that the second constraint only enforces pure lying strategies incur non-positive reward, nevertheless, it is sufficient since any mixed strategy is a convex combination of pure strategies and its corresponding reward is also a convex combination with the same weights. The final constraint assumes all entries of $\mathbf{R}\mathbf{d}$ are negative, thus making sure any report strategy without observing incurs a non-positive reward.
Theorem 3.2 (Optimal cost of a two–player peer-prediction mechanism).
@@
The tight result on the objective function is in the spirit of the classical Crémer-McLean mechanism (cremer1988full), which can extract full surplus from the players when type distributions are common knowledge. The conditions are satisfied for “almost all” $\mathbf{B}$ and $\mathbf{d}$.
@@
Continuing the image labeling example from Example 3.1, we show that the optimal mechanism pays both players the observation cost $c=0.1$ in expectation, as a demonstration of Theorem 3.2. The optimal mechanism can be acquired by solving Eq.(2). Here, we first compute the belief matrix and player $2$’s observation distribution.
@@
Solving the linear program would give the following mechanism: both players receive $5/32$ reward if their reports agree, and receive 5/32-5/32 otherwise.
We now show that the mechanism satisfies all constraints and is cost-optimal. First, no-free-lunch is satisfied since when a player reports without observation, no matter what strategy she follows, expected reward is always $0.5\times 5/32-0.5\times 5/32=0$ (since the other player observes both labels with equal probability). When she lies, expected reward is $(\mathbf{B}\mathbf{R}^{\intercal})_{xy}=-0.1$. When she is truthful, she gets $\sum_{x,x^{\prime}}\mathbb{P}(X_{i}=x)\mathbf{B}_{xx^{\prime}}\mathbf{R}_{xx^{\prime}}=0.1$ in reward, exactly equal to her observation cost. The optimal mechanism extracts full surplus from players.
@@
From Lemma 1 of (radanovic2013robust), it is known that for any mechanism with more than two players with a truthful Bayesian Nash Equilibrium, it is possible to construct a 2-player mechanism (with one focal and one reference player) that achieves a truthful Bayesian Nash Equilibrium with the same expected payment. This lemma narrows our attention to reward mechanisms with only two players. In application, the reference player can be randomly picked to avoid collusion.
@@ 3.2. Distributionally Robust Mechanism Design with Inaccurate Knowledge
Now we move on to the scenario where players and the principal only have inaccurate knowledge of the true $p_{Y}$ and $p_{i}$. Suppose they only know that the true distributions players’ observations $\boldsymbol{X}$ belong to some ambiguity sets $p\in\mathcal{P_{\boldsymbol{X}}}$. From the principal’s perspective, the challenge would be to design a distributionally robust mechanism, such that for any possible realizations within the ambiguity set, the truthfulness constraint would still be met. Its objective now becomes minimizing the expected payment in the worst case.
The notion of distributionally robust mechanisms has been studied in (koccyiugit2020distributionally), which primarily focus on auctions. We focus on the general mechanism design problem, and try to obtain mechanisms that achieve certain guarantees under the worst case, but we don’t pursue exact worst-case optimality. This greatly reduces computational complexity, and turns out to be sufficient for subsequent adaptive mechanism design. From (2) we know the mechanism is evaluated under prior and posterior distribution of the reference player’s observation. With ambiguity set $\mathcal{P_{\boldsymbol{X}}}$, the mechanism design problem can now be formulated as the following distributionally robust optimization problem:
(3) $\begin{split}\min_{R_{i}}\quad&\sup_{p\in\mathcal{P}}\mathbb{E}_{p}[R_{i}(X_{i},X_{j})]\\ \textrm{s.t.}\quad&\inf_{p\in\mathcal{P}_{\boldsymbol{X}}}\mathbb{E}_{p}[R_{i}(X_{i},X_{j})\mid X_{i}]\geq c\\ \quad&\sup_{p\in\mathcal{P}_{\boldsymbol{X}}}\mathbb{E}_{p}[R_{i}(Z_{i},X_{j})\mid X_{i}]\leq c,\quad\forall Z_{i}\neq X_{i}\\ \quad&\sup_{p\in\mathcal{P}_{\boldsymbol{X}}}\mathbb{E}_{p}[R_{i}(Z_{i},X_{j})]\leq 0,\quad\forall Z_{i}\perp\!\!\!\perp X_{i},X_{j}\end{split}$
The constraints are formulated as the three properties are guaranteed even in the worst case. Typically, the analysis revolves around acquiring tight bounds on the objective under different structures of ambiguity set. In the most general case when $\mathcal{P}$ is the set of all distributions on $\boldsymbol{X}$, we would have a mechanism that achieves dominant strategy incentive compatibility (DSIC, in other words, we have a strategyproof mechanism). For our purposes, we don’t need to achieve exact worst-case optimal result on the objective function. Instead, we take an agnostic approach where we don’t require any structure of ambiguity set. All we need to know is the diameter and location of the ambiguity set.
We begin with a sensitivity analysis of (1) with respect to shifts in probability law. Assume the principal obtain a design by solving (1) according to an erroneous probability law $p$, but the true law is $p^{*}$. According to Theorem 3.2, at optimality, we have a binding constraint in (1). Therefore, any slight deviation from $p$ would lead to the potential violation of constraints. To hedge against violations, following the idea of (zhao2024takes), the principal could add safety margins on the constraints. For example, instead of only requiring the expected reward of truthful behaviors to be greater than $c$, the principal could let it be no less than $c+\delta$, where $\delta>0$ is the margin width. In this case, even if the expected reward of truthful behaviors might decrease under $p^{\prime}$, as long as the decrease is no more than $\delta$, the individual rationality property is still preserved. Under this idea, we look at a variant of the mechanism design problem with margin $\delta$:
@@
There are two curious features to this problem variant. First, the margin $\delta$ added to the constraints protects the principal against inaccurate knowledge at an additional cost of at least $\delta$, since the expected payment under each possible observation is at least $c+\delta$. This is a lower bound on the cost of robustness. Increasing $\delta$ means the mechanism from (4) is robust for a higher degree of inaccuracies, but it would also cost more. Pursuing optimality, the principal want to find the lowest $\delta$ just enough to guarantee constraints are still satisfied under $p^{*}$. It would be crucial to understand the connection between the degree of misspecification and the minimal required margin $\delta$. Second, the objective function alters from minimizing expected payment to minimizing worst-case payment. The reason for this change is to limit the sensitivity of expected payment to worst case probability law deviation from $p^{*}$ to $p$. Following the “compactness” criteria discussed in (zhao2024takes), the outcome incurring the highest absolute payment has the highest sensitivity to probability deviation, hence a large $\delta/\kappa$ ratio would ensure a high robustness to such deviations. Therefore, for a fixed $\delta$, we would like to lower $\kappa$ as much as possible.
@@
Let $p$ be the distribution used in designing the margin-$\delta$ mechanism $R_{i}$ in (4) and let $p^{*}$ be the true distribution. Let $p(x_{j}\mid x_{i})$ be the induced prior or posterior distribution of the reference player’s observation $x_{j}$, conditional on the focal player’s observation $x_{i}\in\mathcal{Y}\cup\{\varnothing\}$.
@@
Corollary 3.6 primarily provides the minimum robustness achieved by solving (4). Notice that increasing $\delta$ provides more robustness, but comes at increased cost. Also, the marginal robustness from increasing $\delta$ would diminish: as $\delta\to\infty$, the lower bound is at most $(1-\gamma)/\|\mathbf{B}^{-1}\|((1+\gamma)|\mathcal{Y}|+2)$. When $\delta\to 0$, the robustness provided scales linearly with $\delta$. In addition, from the lower bound on the worst case payments, we have an upper bound $1/2$ on the maximum robustness that can be provided from our scheme. This result does not contradict the impossibility result (Theorem 1) shown in (radanovic2013robust), which proves that no mechanism can guarantee truthfulness for all distributions.
@@
with design distribution $p$ and ambiguity parameter $\eta$. If $\eta\leq\tilde{\eta}=(1-\gamma)/2\|\mathbf{B}^{-1}\|_{2}((1+\gamma)|\mathcal{Y}|+2)$, there exists a mechanism $R_{i}$ such that player $i$ would act truthfully when her belief belongs to this ambiguity set.
@@
Theorem 3.7 is essentially a combination of Theorem 3.4 and 3.5. Notice that the additional cost takes the format of $c\cdot C_{1}\eta/(1-C_{2}\eta)$. This means when $\eta\to 0$, the additional cost of robustness is roughly $O(\eta)$. Also, this theorem holds only when the ambiguity level is lower than a constant threshold $\tilde{\eta}$. This means there is only so much our robustness scheme can do: if the principal starts with no knowledge and the ambiguity level is too high, it is impossible for our robustness mechanism to guarantee truthfulness. We need certain warm-start procedure (for example, an oracle’s opinion) to lower uncertainty under the threshold first.
@@
We follow the same example as in Example 3.1 and 3.3. Now we compare the simple mechanism that pays $1$ on agreement and the optimal mechanism that pays $5/32$. Although the simple mechanism is suboptimal, it is robust to misspecification of players’ skills.
For example, suppose that the players’ true observation accuracy is $0.8$ instead of $0.9$. With the same procedure in Example 3.1, one can show that the simple mechanism still guarantees truthfulness (truthful ($0.26$) ¿ lazy (0) ¿ lying (0.46-0.46). On the other hand, the previously optimal mechanism breaks down (truthfulness gives $9/160<c=0.1$, so players have no incentives in participation). In fact, one can show that as long as both players’ accuracies are the same and stay within the range $[0.66,1]$, the simple mechanism always guarantees truthfulness. The lower bound $(10+\sqrt{10})/20\approx 0.66$ is when the truthful strategy’s expected reward falls to $0.1$. This property holds true even if players know the actual skill level, while the principal does not have that information. In a word, the additional payment in the simple mechanism serves as insurance against ambiguity.

### v2

@@ 2. Problem Formulation
  • reports a fixed permutation of her observation, i.e. for each agent $i$ there exists a bijection $\pi_{i}:\mathcal{Y}\!\to\!\mathcal{Y}$ such that $Z_{it}=\pi_{i}(X_{it})$ .
Proposition 2.1 is derived from Blackwell’s informativeness theorem (Blackwell, 1953). Each round, the true label, observations, reports, and decisions form a Markov chain: $Y_{t}\to\boldsymbol{X}_{t}\to\boldsymbol{Z}_{t}\to A_{t}$. An intuition is that optimal decision-making requires maximal information from upstream. Due to the data processing inequality (Cover, 1999), information never increases going downstream; therefore, the best approach is to preserve as much information as possible at each link. Truthful reporting preserves full information at the link $\boldsymbol{X}_{t}\to\boldsymbol{Z}_{t}$, which allows for optimal subsequent decisions. Any lies from agents erode information. Lazy behavior also produces less information than observing. Aside from truthfulness, an alternative case that preserves full information is when agents use a permutation reporting strategy. However, such a case is unrealistic in practical settings, as the principal would need to know each agent’s permutation rule to reverse the encoding and uncover the true observation. Therefore, this proposition essentially shows that eliciting truthfulness is the practical way to achieve maximal performance.
Proposition 2.1 also shows that truthfulness is necessary not only for data quality but also cost-optimality. This is because the design of cost-optimal mechanisms is itself a decision-making task and Proposition 2.1 applies.
@@ 3. Mechanism Design without Common Knowledge
In our work, a central relaxation of modeling is that we don’t assume prior distribution of labels $p_{Y}$ or agents’ skills $p_{i}$ are known by agents or the principal. The principal’s attempts to maintain truthfulness with unknown or inaccurate estimation of such knowledge. In this section, we focus on distributionally robust mechanisms, which aim to incentivize truthful behavior under knowledge ambiguity.
@@ 3.1. Optimal Single-round Mechanism Design
We begin with the analysis of optimal mechanism design with known $p_{Y}$ and $p_{i}$ within a single round. When there are no true labels available, we apply the principles of peer prediction, which is to use other agents’ report to verify a focal agent’s report. The delicacy lies in the careful design of the reward mechanism to ensure that truthfulness is a Nash Equilibrium.
We start with the two-agent mechanism. With a focal agent $i$ and a reference agent $j$, the optimal two-agent mechanism design problem could be formulated as a linear programming problem. The objective is to minimize expected reward to agents, and the constraints are the desired properties of the mechanism. (we hide subscript $t$ for simplicity.)
@@
The expectation is taken under the joint probability law $p_{\boldsymbol{X}}$ induced by $p_{Y},p_{i},$ and $p_{j}$. In fact, $p_{\boldsymbol{X}}$ uniquely determines the mechanism design problem. The first constraint is the individual rationality property, meaning when the agent obtains observations and then reports honestly, she would get a non-negative reward. The second one states the truthfulness property, where the agent receives a non-positive reward when she lies. The final constraint implements the no-free-lunch property, meaning the agent cannot get a positive utility when she is lazy. We introduce the individual rationality and the no-free-lunch constraint to prevent arbitrary decrease of the objective function by applying an affine transformation to the reward mechanisms. (For risk-neutral agents, affine transformations on reward do not affect utility ordering and strategic behavior.)
@@
Suppose there are two types of images $\mathcal{Y}=\{\texttt{Cat},\texttt{Tiger}\}$, abbreviated with $C$ and $T$ respectively. We further assume that the prior distribution of the image types is balanced, i.e., $p_{Y}(C)=p_{Y}(T)=0.5$. For each image with an unknown true label $Y\in\mathcal{Y}$, the principal would like to let two agents $1,2$ individually observe it, and truthfully report their observations $X_{1},X_{2}$ to label that image. Assume that both agents are $90\%$ accurate: $p_{i}(C\mid C)=p_{i}(T\mid T)=0.9$, and $p_{i}(T\mid C)=p_{i}(C\mid T)=0.1$.
The principal designs the reward mechanism as follows: both agents receive $1$ reward if their reports $Z_{1},Z_{2}$ agree with each other, and receive 1-1 otherwise, i.e. $R_{\{1,2\}}(Z_{1},Z_{2})=2\cdot\mathbf{1}_{[Z_{1}=Z_{2}]}-1$. and assume that observation incurs cost $c=0.1$. Now we assume that agent $2$ observes and report honestly, and analyze the incentive of agent $1$.
Suppose agent $1$ observes Cat, Bayes’ formula gives that $\mathbb{P}(X_{2}=C\mid X_{1}=C)=0.82$ and $\mathbb{P}(X_{2}=C\mid X_{1}=C)=0.18$. On the other hand, if agent $1$ does not pay the effort to toss the coin, then her Bayesian belief on agent $2$’s observation (and report) is $P(X_{2}=C)=P(X_{2}=T)=0.5$. She can then work out expected reward under truthful, lying, and lazy strategies: truthful ($0.54$) ¿ lazy (0) ¿ lying (0.74-0.74). Hence truthful behavior is desired. In fact, this simple mechanism is a feasible solution to Eq.(1). Intuitively, after the focal agent’s observation, her posterior probability on the other agent’s observing the same label is higher than observing a different label. Therefore, it is preferable to report whatever you observe in the first place. Such a mechanism is called peer prediction (Miller et al., 2005), originated the fact that rational agents always tries to predict their peers’ observations before action.
Define the belief matrix $\mathbf{B}$ where $\mathbf{B}_{xx^{\prime}}=\mathbb{P}(X_{j}=x^{\prime}\mid X_{i}=x)$, and reward matrix $\mathbf{R}$ where $\mathbf{R}_{xx^{\prime}}=R_{i}(x,x^{\prime})$. We also let $\mathbf{d}$ be a column vector display of $j$’s observation distribution: $\mathbf{d}_{x}=\mathbb{P}(X_{j}=x)$. Then we can reformulate (1) into the following equivalent problem:
@@
Note that the second constraint only enforces pure lying strategies to incur non-positive reward, nevertheless, it is sufficient since any mixed strategy is a convex combination of pure strategies and its corresponding reward is also a convex combination with the same weights. The final constraint assumes all entries of $\mathbf{R}\mathbf{d}$ are negative, thus making sure any report strategy without observing incurs a non-positive reward.
Theorem 3.2 (Optimal cost of a two–agent peer-prediction mechanism).
@@
The tight result on the objective function is in the spirit of the classical Crémer-McLean mechanism (Crémer and McLean, 1988), which can extract full surplus from the agents when type distributions are common knowledge. The conditions are satisfied for “almost all” $\mathbf{B}$ and $\mathbf{d}$.
@@
Continuing the image labeling example from Example 3.1, we show that the optimal mechanism pays both agents the observation cost $c=0.1$ in expectation, as a demonstration of Theorem 3.2. The optimal mechanism can be acquired by solving Eq.(2). Here, we first compute the belief matrix and agent $2$’s observation distribution.
@@
Solving the linear program would give the following mechanism: both agents receive $5/32$ reward if their reports agree, and receive 5/32-5/32 otherwise.
We now show that the mechanism satisfies all constraints and is cost-optimal. First, no-free-lunch is satisfied since when a agent reports without observation, no matter what strategy she follows, expected reward is always $0.5\times 5/32-0.5\times 5/32=0$ (since the other agent observes both labels with equal probability). When she lies, expected reward is $(\mathbf{B}\mathbf{R}^{\intercal})_{xy}=-0.1$. When she is truthful, she gets $\sum_{x,x^{\prime}}\mathbb{P}(X_{i}=x)\mathbf{B}_{xx^{\prime}}\mathbf{R}_{xx^{\prime}}=0.1$ in reward, exactly equal to her observation cost. The optimal mechanism extracts full surplus from agents.
@@
From Lemma 1 of (Radanovic and Faltings, 2013), it is known that for any mechanism with more than two agents with a truthful Bayesian Nash Equilibrium, it is possible to construct a 2-agent mechanism (with one focal and one reference agent) that achieves a truthful Bayesian Nash Equilibrium with the same expected payment. This lemma narrows our attention to reward mechanisms with only two agents. In application, the reference agent can be randomly picked to avoid collusion.
@@ 3.2. Distributionally Robust Mechanism Design with Inaccurate Knowledge
Now we move on to the scenario where agents and the principal only have inaccurate knowledge of $p_{\boldsymbol{X}}$. Suppose they only know that the true distributions agents’ observations $\boldsymbol{X}$ belong to some ambiguity sets $p\in\mathcal{P_{\boldsymbol{X}}}$. From the principal’s perspective, the challenge would be to design a distributionally robust mechanism, such that for any possible realizations within the ambiguity set, the truthfulness constraint would still be met. Its objective now becomes minimizing the expected payment in the worst case.
The notion of (distributionally) robust mechanisms has been studied in (Bergemann and Morris, 2005; Koçyiğit et al., 2020). We focus on a specific family that is cheap enough and guarantees truthfulness under ambiguity, but we don’t pursue exact worst-case optimality. This greatly reduces computational complexity, and turns out to be sufficient for subsequent adaptive mechanism design.
We begin with a sensitivity analysis of (1) with respect to shifts in probability law. Assume the principal obtain a design by solving (1) according to an erroneous probability law $p$, but the true law is $p^{*}$. According to Theorem 3.2, at optimality, we have a binding constraint in (1). Therefore, any slight deviation from $p$ would lead to the potential violation of constraints. To hedge against violations, following the idea of (Zhao et al., 2024), the principal could add safety margins on the constraints. Instead of only requiring the expected reward of truthful behaviors to be greater than $c$, the principal could let it be no less than $c+\delta$, where $\delta>0$ is the margin width. In this case, even if the expected reward of truthful behaviors might decrease under $p^{\prime}$, as long as the decrease is no more than $\delta$, the individual rationality property is still preserved. Under this idea, we look at a variant of the mechanism design problem with margin $\delta$:
@@
There are two curious features to this problem variant. First, the margin $\delta$ added to the constraints protects the principal against inaccurate knowledge at an additional cost of at least $\delta$, since the expected payment under each possible observation is at least $c+\delta$. This is a lower bound on the cost of robustness. Increasing $\delta$ means the mechanism from (3) is robust for a higher degree of inaccuracies, but it would also cost more. Pursuing optimality, the principal want to find the lowest $\delta$ just enough to guarantee constraints are still satisfied under $p^{*}$. It would be crucial to understand the connection between the degree of misspecification and the minimal required margin $\delta$. Second, the objective function alters from minimizing expected payment to minimizing worst-case payment. The reason for this change is to limit the sensitivity of expected payment to worst case probability law deviation from $p^{*}$ to $p$. Following the “compactness” criteria discussed in (Zhao et al., 2024), the outcome incurring the highest absolute payment has the highest sensitivity to probability deviation, hence a large $\delta/\kappa$ ratio would ensure a high robustness to such deviations. Therefore, for a fixed $\delta$, we would like to lower $\kappa$ as much as possible.
@@
Let $p$ be the distribution used in designing the margin-$\delta$ mechanism $R_{i}$ in (3) and let $p^{*}$ be the true distribution. Let $p(x_{j}\mid x_{i})$ be the induced prior or posterior distribution of the reference agent’s observation $x_{j}$, conditional on the focal agent’s observation $x_{i}\in\mathcal{Y}\cup\{\varnothing\}$.
@@
Corollary 3.6 primarily provides the minimum robustness achieved by solving (3). Notice that increasing $\delta$ provides more robustness, but comes at increased cost. Also, the marginal robustness from increasing $\delta$ would diminish: as $\delta\to\infty$, the lower bound is at most $(1-\gamma)/\|\mathbf{B}^{-1}\|((1+\gamma)|\mathcal{Y}|+2)$. When $\delta\to 0$, the robustness provided scales linearly with $\delta$. In addition, from the lower bound on the worst case payments, we have an upper bound $1/2$ on the maximum robustness that can be provided from our scheme. This result does not contradict the impossibility result (Theorem 1) shown in (Radanovic and Faltings, 2013), which proves that no mechanism can guarantee truthfulness for all distributions.
@@
with design distribution $p$ and ambiguity parameter $\eta$. If $\eta\leq\tilde{\eta}=(1-\gamma)/2\|\mathbf{B}^{-1}\|_{2}((1+\gamma)|\mathcal{Y}|+2)$, there exists a mechanism $R_{i}$ such that agent $i$ would act truthfully when her belief belongs to this ambiguity set.
@@
Theorem 3.7 is essentially a combination of Theorem 3.4 and 3.5. Notice that the additional cost takes the format of $c\cdot C_{1}\eta/(1-C_{2}\eta)$. This means when $\eta\to 0$, the additional cost of robustness is roughly $O(\eta)$, and we establish a linear relation between robustness and additional cost.
@@
We follow the same example as in Example 3.1 and 3.3. Now we compare the simple mechanism that pays $1$ on agreement and the optimal mechanism that pays $5/32$. Although the simple mechanism is suboptimal, it is robust to misspecification of agents’ skills.
For example, suppose that the agents’ true observation accuracy is $0.8$ instead of $0.9$. With the same procedure in Example 3.1, one can show that the simple mechanism still guarantees truthfulness (truthful ($0.26$) ¿ lazy (0) ¿ lying (0.46-0.46). On the other hand, the previously optimal mechanism breaks down (truthfulness gives $9/160<c=0.1$, so agents have no incentives in participation). In fact, one can show that as long as both agents’ accuracies are the same and stay within the range $[0.66,1]$, the simple mechanism always guarantees truthfulness. The lower bound $(10+\sqrt{10})/20\approx 0.66$ is when the truthful strategy’s expected reward falls to $0.1$. This property holds true even if agents know the actual skill level, while the principal does not have that information. In a word, the additional payment in the simple mechanism serves as insurance against ambiguity.